Find the mass in grams of 5.50 x 10^23 molecules of F2
1
Expert's answer
2019-04-18T05:05:36-0400

According to the moles-mass equation:

m = n × Mr = (N/NA) × Mr,

where n - number of moles, Mr - molecular weight, N - number of molecules, NA - Avogadro's number.

Here, N = 5.50 × 1023, Mr = 38 g/mol, NA = 6.022 × 1023.

As a result, the mass of F2 equals:

m = (5.50 × 1023 / 6.022 × 1023) × 38 g/mol = 0.913 mol × 38 g/mol = 34.7 g.

Answer: 34.7g of F2.